# EndoDDC: Learning Sparse to Dense Reconstruction for Endoscopic Robotic Navigation via Diffusion Depth Completion

### v1

@@ IV Experiments
RMSE ($m$) ↓ MAE ($m$) ↓ REL ($m$) ↓ $\delta$ RMSE ($m$) ↓ MAE ($m$) ↓ REL ($m$) ↓ $\delta$
@@ IV-A Datasets and Evaluation Metrics
Evaluation Metrics. We evaluate under 4 standard metrics [24, 23, 12]: root mean squared error RMSE, mean absolute error MAE, mean absolute relative error REL in millimeters ($m$), and $\delta$ accuracy, which measures the percentage of predicted pixels $\hat{d}$ that satisfy $\max(\frac{\hat{d}}{d},\frac{d}{\hat{d}})<1.25^{k}$, where $k=1$. This metric provides a robust measure of the proportion of depth estimates within a clinically acceptable error margin.
@@ IV-B Implementation Details
RMSE ($m$) ↓ REL ($m$) ↓ RMSE ($m$) ↓ REL ($m$) ↓ RMSE ($m$) ↓ REL ($m$) ↓ RMSE ($m$) ↓ REL ($m$) ↓
@@
RMSE ($m$) ↓ MAE ($m$) ↓ REL ($m$) ↓ $\delta$ RMSE ($m$) ↓ MAE ($m$) ↓ REL ($m$) ↓ $\delta$
@@ IV-E Ablation Study
In the ablation study, we compare our method with the baseline [47, 37] and the DDI [49] method for iterative update. For fair comparison, we report the results of the baseline without the Depth Diffusion and with $D$. To enable fair comparison, we fine-tune the baseline method on the surgical datasets [2, 8]. Specifically, we initialized the baseline models with weights pre-trained on natural image datasets and subsequently trained them on C3VD and StereoMIS using the same optimizer and learning rate as EndoDDC. This process ensures that the performance gap is not due to domain shift between natural and surgical environments, but rather stems from the architectural advantages of our proposed diffusion-based refinement. We also compare our full method with the version using depth diffusion without the Depth Grad Fusion feature as diffusion condition, and the version of depth diffusion with depth gradient condition but without the initial depth. The former performs diffusion without guidance of depth and gradient features, while the latter does not incorporate initial depth information during denoising, instead initiating the diffusion process directly from random noise. The results are presented in Table III. Compared with others, our method shows dramatic improvement across all metrics. This result shows that our proposed Depth Grad Fusion module (for diffusion conditioning) and Depth Diffusion module with initial depth significantly optimize the depth map, leading to state-of-the-art performance by effectively modeling the data distribution and reducing residual errors.

### v2

@@ IV Experiments
RMSE (mm) ↓ MAE (mm) ↓ REL (mm) ↓ $\delta$ RMSE (mm) ↓ MAE (mm) ↓ REL (mm) ↓ $\delta$
@@ IV-A Datasets and Evaluation Metrics
Evaluation Metrics. We evaluate under 4 standard metrics [24, 23, 12]: root mean squared error RMSE, mean absolute error MAE, mean absolute relative error REL in millimeters (mm), and $\delta$ accuracy, which measures the percentage of predicted pixels $\hat{d}$ that satisfy $\max(\frac{\hat{d}}{d},\frac{d}{\hat{d}})<1.25^{k}$, where $k=1$. This metric provides a robust measure of the proportion of depth estimates within a clinically acceptable error margin.
@@ IV-B Implementation Details
RMSE (mm) ↓ REL (mm) ↓ RMSE (mm) ↓ REL (mm) ↓ RMSE (mm) ↓ REL (mm) ↓ RMSE (mm) ↓ REL (mm) ↓
@@
RMSE (mm) ↓ MAE (mm) ↓ REL (mm) ↓ $\delta$ RMSE (mm) ↓ MAE (mm) ↓ REL (mm) ↓ $\delta$
@@ IV-E Ablation Study
In the ablation study, we compare our method with the baseline [47, 37] and the DDI [49] method for iterative update. For fair comparison, we report the results of the baseline without the Depth Diffusion and with DD. To enable fair comparison, we fine-tune the baseline method on the surgical datasets [2, 8]. Specifically, we initialized the baseline models with weights pre-trained on natural image datasets and subsequently trained them on C3VD and StereoMIS using the same optimizer and learning rate as EndoDDC. This process ensures that the performance gap is not due to domain shift between natural and surgical environments, but rather stems from the architectural advantages of our proposed diffusion-based refinement. We also compare our full method with the version using depth diffusion without the Depth Grad Fusion feature as diffusion condition, and the version of depth diffusion with depth gradient condition but without the initial depth. The former performs diffusion without guidance of depth and gradient features, while the latter does not incorporate initial depth information during denoising, instead initiating the diffusion process directly from random noise. The results are presented in Table III. Compared with others, our method shows dramatic improvement across all metrics. This result shows that our proposed Depth Grad Fusion module (for diffusion conditioning) and Depth Diffusion module with initial depth significantly optimize the depth map, leading to state-of-the-art performance by effectively modeling the data distribution and reducing residual errors.